Let {an}∞n=1 be a convergent sequence. Prove that {an}∞n=1 satisfies Cauchy’s criterion


1
Expert's answer
2021-02-24T06:39:38-0500

"\mathrm{Given:\;\{a_n\}_{n=1}^{\infty}\;is\;a\;convergent\;sequence.}"


"\mathrm{Prove:\{a_n\}_{n=1}^{\infty}\;satisfies\;Cauchy's\;criterion.}"


"\mathrm{Proof:\;Let\;\epsilon >0\;be \;given.Since\;\{a_n\}_{n=1}^{\infty}\;is\;a\;convergent\;sequence.\;So,}"

"\mathrm{choose\;N\;such\;that\;if\;n>N,\;we\;have\;|a_n-\alpha |<\dfrac{\epsilon}{2}\;.}"


"\mathrm{Then\;if\;m,n>N,we\;have\;|a_n-a_m|=|a_n-\alpha+\alpha-a_m|}"

"\mathrm{\leq |a_n-\alpha|+|a_m-\alpha|}"

"\mathrm{< \dfrac{\epsilon}{2}+\dfrac{\epsilon}{2}}"

"\mathrm{=\epsilon}"


"\mathrm{\therefore \{a_n\}_{n=1}^{\infty}\;is\;a\;cauchy\;sequence.}"



"\mathrm{}"